# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v8

@@ 1 Introduction
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. Kim et al. (2018b) discussed the problem of performing LR training in an encrypted environment. They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function. In the iDASH 2017 competition, Bonte and Vercauteren Bonte & Vercauteren (2018), Kim et al. Kim et al. (2018a), Chen et al. Chen et al. (2018), and Crawford et al. Crawford et al. (2018) all investigated the same problem that Kim et al. Kim et al. (2018b) studied. In the iDASH competition of 2018, Kim et al. Kim et al. (2019b) and Blatt et al. Blatt et al. (2019) further worked on it for an efficient packing and semi-parallel algorithm. There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a). Bonte and Vercauteren Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. Kim et al. (2018a) for efficient homomorphic computation.
@@
    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms in most cases demonstrate high performance in terms of convergence speed.
@@ 2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (Gentry, 2009). FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown. Cheon et al. Cheon et al. (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a; b; Han et al., 2019).
@@ 2.2 Database Encoding Method
Kim et al. Kim et al. (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
@@
Han et al. Han et al. (2019) mentioned several basic but important operations used by Kim et al. Kim et al. (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
@@ 3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning and Lindsay Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
@@ 3.1 Simplified Fixed Hessian
Bonte and Vercauteren Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
@@
The SFH method proposed by Bonte and Vercauteren Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.
@@ 3.2 Quadratic Gradient Definition
As long as $\tilde{B}$ satisfies the convergence condition of the above fixed Hessian method, $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound. Since we already assume that $\bar{H}\leq H$, it will suffice to show that $\tilde{B}\leq\bar{H}$. We prove $\tilde{B}\leq\bar{H}$ in a similar way that (Bonte & Vercauteren, 2018) did.
@@
Given such a $\tilde{B}$ above, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $\eta$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$, and $\eta$ should be always no less than 1 and decrease to 1 in a limited number of iteration steps. Note that $G$ is still a column vector of the same size as the gradient $\bm{g}$. To maximize the function $F(\mathbf{x})$, we can use the iterative formulas: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+\eta\cdot G$, just like the naive gradient.
@@
Böhning and Lindsay Böhning & Lindsay (1988) did not proposed a systemic way to find or build the constant Hessian approximate for their fixed Hessian method. Perhaps there are no such matrices for most objective functions. We mention here that our quadratic gradient doesn’t need a constant (fixed) Hessian replacement to construct and that we can use the original Hessian itself to build our varying diaginal matrix $\tilde{B}$ for the quadratic gradient. Because of this loose condition, we propose a system way satisfying sufficient condition to find if there is a fixed Hessian matrix for efficient calculation. That is, we first build $\tilde{B}$ from the Hessian of the given objective function and then check if for every its diaginal element $\tilde{B}_{kk}$ there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or constant minimum $|\tilde{B}_{kk}|$). If each diagnal element $|\tilde{B}_{kk}|$ all had their minimum $|\tilde{M}_{kk}|$, then the constant diagnial matrix with corresponding diagnal elements $|\tilde{M}_{kk}|$ would satisfy the convengence condiion for SFH method. Finally, we should also evaluate the performance of such fixed Hessian replacement since it might not be a $good$ bound. This searching method can be employed to any optimization functions. For example, we can build $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ for the logistic regressian and attempt to maximise $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The resulting constant diagnoal matrix is the same as the one built from the $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction.
Böhning and Lindsay Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient. Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation. Specifically, we first construct $\tilde{B}$ from the Hessian of the given objective function and then check if, for each diagonal element $\tilde{B}_{kk}$, there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or a constant minimum $|\tilde{B}_{kk}|$). If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method. Finally, we should evaluate the performance of such a fixed Hessian replacement, as it might not always provide a good bound. This search method can be applied to any optimization function to identify a fixed Hessian. However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
Our method can be employed to identify a fixed Hessian for binary logistic regression. For instance, we can construct $|\tilde{B}_{kk}|$ from the Hessian $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ and then try to maximize $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The resulting constant diagonal matrix is equivalent to the one obtained from $-\frac{1}{4}X^{\top}X$ using the quadratic-gradient construction. Therefore, $-\frac{1}{4}X^{\top}X$ could serve as the fixed Hessian for the fixed Hessian method, precisely as suggested by Böhning and Lindsay Böhning & Lindsay (1988).
@@ 3.3 Quadratic Gradient Algorithms
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The enhanced Adagrad algorithm use $N_{t}=1+0.01$ as the learning rate while the enhanced Adam algorithms adopt the following learning rate setting: $\alpha=1+0.001,$ $\beta_{1}=0.9,$ and $\beta_{2}=0.999.$ In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithms outperform the original NAG method in terms of convergence rate. Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases. The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work. One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
@@ 4 Secure Training
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned drawbacks.

### v9

@@ 1 Introduction
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. (2018b) discussed the problem of performing LR training in an encrypted environment. They employed full-batch gradient descent during the training process and utilized the least-squares method to approximate the sigmoid function. In the 2017 iDASH competition, Bonte & Vercauteren (2018), Kim et al. (2018a), Chen et al. (2018), and Crawford et al. (2018) all addressed the same problem explored by Kim et al. (2018b). In the 2018 iDASH competition, Kim et al. (2019b) and Blatt et al. (2019) further developed the problem, focusing on efficient packing and a semi-parallel algorithm. There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a). Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. (2018a) for efficient homomorphic computation.
@@
    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms all converge and, in most cases, demonstrate strong performance in terms of convergence speed.
@@ 2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (Gentry, 2009). FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown. Cheon et al. (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a; b; Han et al., 2019).
@@ 2.2 Database Encoding Method
Kim et al. (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose a database contains a training dataset $Z$ comprising $n$ samples and $(1+d)$ covariates. The dataset matrix $Z$ is packed into a single ciphertext in a row-by-row manner.
@@
Han et al. (2019) mentioned several basic but important operations used by Kim et al. (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
@@ 3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
@@ 3.1 Simplified Fixed Hessian
Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
@@
The SFH method proposed by Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.
@@ 3.2 Quadratic Gradient Definition
As long as $\tilde{B}$ satisfies the convergence condition of the aforementioned fixed Hessian method, namely $\tilde{B}\leq H$, we can use this approximation $\tilde{B}$ of the Hessian matrix as a lower bound. Given that we have already assumed $\bar{H}\leq H$, it suffices to demonstrate that $\tilde{B}\leq\bar{H}$. We establish $\tilde{B}\leq\bar{H}$ in a manner similar to that employed by Bonte & Vercauteren (2018).
@@
Given the previously defined $\tilde{B}$, we define the quadratic gradient as $G=\bar{B}\cdot\bm{g}$ with a new learning rate $N_{t}$, where $\bar{B}$ is a diagonal matrix with diagonal entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. The learning rate $N_{t}$ should always be no less than 1 and is designed to decrease to 1 over a limited number of iterations. It is important to note that $G$ remains a column vector of the same dimension as the gradient $\bm{g}$. To maximize the function $F(\mathbf{x})$, we can employ the iterative formula: $\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot G$, similar to the naive gradient approach.
@@
Böhning & Lindsay (1988) did not propose a systematic method for determining or constructing a constant Hessian approximation for their fixed Hessian method. This may be due to the absence of such matrices for most objective functions. We note that our quadratic gradient approach does not require a constant (fixed) Hessian replacement and instead allows for the use of the original Hessian itself to construct our varying diagonal matrix $\tilde{B}$ for the quadratic gradient. Under this less restrictive condition, we present a methodical approach to evaluate whether a fixed Hessian matrix can be employed for efficient computation. Specifically, we first construct $\tilde{B}$ from the Hessian of the given objective function and then check if, for each diagonal element $\tilde{B}_{kk}$, there is a constant maximum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$ (or a constant minimum $|\tilde{B}_{kk}|$). If each diagonal element $|\tilde{B}_{kk}|$ achieves its minimum $|\tilde{M}_{kk}|$, then a constant diagonal matrix with the corresponding diagonal elements $|\tilde{M}_{kk}|$ would satisfy the convergence condition for the SFH method. Finally, we should evaluate the performance of such a fixed Hessian replacement, as it might not always provide a good bound. This search method can be applied to any optimization function to identify a fixed Hessian. However, the absence of a constant Hessian replacement via this method does not necessarily imply that such an approximation does not exist.
@@ 3.3 Quadratic Gradient Algorithms
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function $l(\bm{\beta})$, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (edin), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work by Kim et al. (2018a), the enhanced NAG algorithm employs the learning rate $1+\frac{10}{1+t}$, consistent with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. The enhanced Adagrad algorithm utilizes $N_{t}=1+0.01$ as the learning rate, while the enhanced Adam algorithm adopts the following learning rate settings: $\alpha=1+0.001$, $\beta_{1}=0.9$, and $\beta_{2}=0.999$. Refer to (Kingma & Ba, 2014) for further details regarding these parameters. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the function $l(\bm{\beta})$.
From the empirical results presented in Figures 12, and 3, we conclude that all the enhanced algorithms converge, and that the enhanced NAG algorithm consistently outperforms the original NAG method in terms of convergence rate. Although the enhanced Adagrad and Adam methods do not achieve faster convergence than their original counterparts on the iDASH genomic dataset, they demonstrate clear advantages in all other cases. The reason for this discrepancy is beyond the scope of this paper and will be addressed as part of future work. One possible explanation for the superior performance of the enhanced algorithms is that the quadratic gradient integrates curvature information into first-order gradient methods.
@@ 4 Secure Training
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
In a privacy-preserving setting, when compared to the NAG method, the primary limitation of the Enhanced NAG method is that it requires one additional ciphertext multiplication to construct the quadratic gradient. In addition, the data owner needs to upload one more ciphertext encrypting the matrix $\bar{B}$. However, the enhanced algorithm converges faster, and we believe it can compensate for the mentioned shortcomings.